# Privacy Guarantees in Posterior Sampling under Contamination

### v2

@@ 4.1 Estimation Procedure
There are a few places where the use of $\theta^{*}$ is preferred. For the simulations, we can use $\theta^{*}$ directly, but when $\theta^{*}$ is unknown, an estimation $\hat{\theta}^{*}$ would suffice.
Remark 15.
In order for Assumption 4 to be satisfied for regression settings, one needs to impose a constraint on the size of covariates. In theory, we may allow the size of $\bm{W}$ to grow with $n$, for simplicity, we assume in all the simulations that $\|\bm{W}\|_{\infty}\leq 1$.
input : Constants $C_{\phi}^{(1)},C_{\phi}^{(2)},C_{p}$; Likelihood function $f$; Contamination density $h$; prior $\pi_{0}$; Data set size $n$; Number of repeats $K$; Number of particles $m$; Estimated true parameter $\hat{\theta}^{*}$;
1 Compute $\phi_{n},p_{n}$ in (21);
2 Set $A_{n}:=B_{\phi_{n}}(\hat{\theta}^{*},\infty)$ to be the $L_{\infty}$-ball around $\hat{\theta}^{*}$, i.e., a hypercube of width $2\phi_{n}$.;
@@ 3 Solve
4 for $k=1,\dots,K$ do
5       if No dataset is present then
6             Generate covariates $\bm{W}:=\{\bm{w}_{1},\dots,\bm{w}_{n}\}$ from a desired distribution, e.g., standard multivariate normal.;
7             Generate true observations $\bm{X}:=\{\bm{x}_{1},\dots,\bm{x}_{n}\}$ according to $\bm{x}_{i}\sim f(\cdot;\theta^{*}|\bm{w}_{i})$;
8
9       Contaminate the observations $\bm{X}$, denoted $\tilde{\bm{X}}$, by replacing each $\bm{x}_{i}$ with probability $p_{n}$ with a draw from $g(\cdot)$;
10       Compute the Laplace approximation of the contaminated posterior distribution $\pi_{n}(\cdot;\tilde{\bm{X}})$ according to (19) ;
11       Sample $\theta_{1},\dots,\theta_{m}$ from the Laplace approximation;
12       Compute normalized importance weights $w_{1},\dots,w_{m}$ with respect to the $\pi_{n}(\cdot;\tilde{\bm{X}})$;
13       Compute $\hat{\delta}_{n}^{(k)}$ in (20);
@@ 14       Set $\mathcal{D}$ to be $\tilde{\bm{X}}$ with the last element discarded;
15       Compute the normalized weight $\tilde{w}_{i}$ with respect to $\pi_{n-1}(\cdot;\mathcal{D})$;
@@ 16       Solve for
$\alpha:=\sup_{\bm{x}}\sum_{i=1}^{m}w_{i}d(\bm{x};\theta_{i}),\quad\beta:=\inf_{\bm{z}}\sum_{i=1}^{m}w_{i}d(\bm{x};\theta_{i})$
@@
17       Set $\hat{\epsilon}_{n}^{(k)}:=\eta_{n}+\log(\alpha)-\log(\beta)$;
18
@@ 17       Set $\hat{\epsilon}_{n}^{(k)}:=\eta_{n}+\log(\alpha)-\log(\beta)$;
output : Estimated $\hat{\epsilon}_{n}^{(k)}$, $\hat{\delta}_{n}^{(k)}$, $k=1,\dots,K$
@@
Algorithm 1 An empirical procedure to estimate $\epsilon$ and $\delta$ for a given set of parameters $A_{n},p_{n}$.
@@ 4.2 Simulation Setups and Results
Let $\bm{X}=\{X_{1},\dots,X_{n}\}\in\mathbb{R}^{n}$ be the set of observations and $\bm{W}\in[-1,1]^{n\times d}$ be the covariate matrix associated with $\bm{X}$, i.e., the $i$-th row of $\bm{W}$ is associated with the $i$-th observation $X_{i}$. By default, we assume the first column of $\bm{W}$ is a vector of $1$. Let $\theta\in\mathbb{R}^{d}$ denote the parameter and the model likelihood is given by a density function $f(X_{i};\theta,\bm{w}_{i})=f(X_{i},\theta^{\top}\bm{w}_{i})$, such that $\int_{\mathbb{R}}f(x;\theta,\bm{w})\mathtt{d}x=1$ for any pair of $\theta,\bm{w}\in\mathbb{R}^{d}$. Let $\theta^{*}$ denote the true parameter value, which in practice may be replaced by an estimation $\hat{\theta}$, e.g., using the maximum a posteriori estimator.
For the cases where contamination is required, we choose a contamination density $g(x;\bm{w})$ and replace each $X_{i}$ with probability $p_{n}$ by a random sample from $g(\cdot;\bm{w}_{i})$. Note that the contamination density is allowed to depend on the covariate $\bm{w}$, see the example below for the linear regression case.
@@
$k_{p_{n}}(x;\theta,\bm{w})=(1-p_{n})f(x;\theta,\bm{w})+p_{n}\,g(x;\bm{w}).$
The prior $\pi_{0}$ on $\theta$ is always set as a zero mean Gaussian prior with the appropriate dimension. The covariance is a diagonal matrix with all diagonal terms equal to $100$.
@@
    Bayesian Linear Regression: ($C_{\phi}^{(1)}=5$)
    The regression model is given by
    $X_{i}=\theta^{*\top}\bm{w}_{i}+e_{i}$
    where $e_{i}\sim\mathcal{N}(0,1)$. The likelihood function $f(X;\theta,\bm{w})$ is given by
    $f(X;\theta,\bm{w})=\frac{1}{2\pi\sigma^{2}}\exp\left(-\frac{(X-\theta^{\top}\bm{w})^{2}}{2}\right).$
    The contamination density is chosen as $g(X_{i};\bm{w}_{i})\sim T_{5}(\theta^{*\top}\bm{w}_{i})$, a shifted Student’s t-distribution with degrees of freedom $5$, given by the following expression
    $g(x;\bm{w})=\frac{\Gamma(3)}{\sqrt{3\pi}\Gamma(\tfrac{5}{2})}\left(1+\frac{(x-\theta^{*\top}\bm{w})^{2}}{5}\right)^{-3}$
    where $\sigma$ is the standard deviation of the model noise for convenience. In this setup, a heavy-tailed contamination is required in order to remove the compactness constraint on the observation space.
@@
    Bayesian Logistic Regression: ($C_{\phi}^{(1)}=10$)
    The regression model is given by $X_{i}\sim\text{Ber}\left((1+e^{-\theta^{*\top}\bm{w}_{i}})^{-1}\right)$ where $X\in\{0,1\}$ with likelihood function
    $f(X;\theta,\bm{w})=\left(1+e^{-\theta^{*\top}\bm{w}_{i}}\right)^{-X}\left(1+e^{\theta^{*\top}\bm{w}_{i}}\right)^{-1+X}.$
    Since the observation space is compact, $\epsilon$ is bounded in this setup. However, as we will see in Section 16, contamination should still be adopted. To contaminate the observations in this case, we replace each observation $X_{i}$ with probability $p_{n}$ by a draw from $\text{Ber}(\tfrac{1}{2})$. Thus the contaminated density is $k_{p_{n}}(X;\theta,\bm{w})=(1-p_{n})f(X;\theta,\bm{w})+0.5p_{n}.$
@@
    Bayesian Cauchy Regression: ($C_{\phi}^{(1)}=8$)
    The regression model is given by $X_{i}=\theta^{*\top}\bm{w}_{i}+e_{i}$ where $e_{i}\sim\text{Cauchy}(0)$ with likelihood function
@@
    The simulations are done without contamination as $d(X;\theta)$ has bounded sensitivity.
Each of the above setups is simulated with parameter dimension $=51$ and the estimated $(\epsilon_{n},\delta_{n})$ results are plotted in Figure 1. The three rows each represent the results for Linear regression, Logistic regression (with contamination) and Cauchy regression respectively.
For each combination of $(C_{\phi}^{(1)},C_{\phi}^{(2)},C_{p})$, the simulation is repeated for $48$ times. For each combination of $(C_{\phi}^{(1)},C_{\phi}^{(2)},C_{p})$, the average $\hat{\epsilon}_{n},\hat{\delta}_{n}$ over all repetitions, and the maximum $\hat{\epsilon}_{n},\hat{\delta}_{n}$ over all repetitions are plotted. The max curves and average curves are distinguished by the point shape, circle for average and triangle for max. The curves are plotted against $n$, the number of data points in the dataset. We see in all cases, the max curve and the average curve shrink at a similar rate.
Figure 1: Plots of estimated $(\epsilon_{n},\delta_{n})$ for three different model setups, with parameter dimension $51$. The average and maximum across all repeats are plotted and distinguished by dot shape. Each colour and line type pair corresponds to the result of one choice of the constant.
(a) Linear regression case
(b) Logistic regression case
(c) Cauchy regression case
(d) Uncontaminated logistic regression
Remark 16 (Contaminated Logistic Regression).
Similar to the Cauchy case, Logistic regression is differentially private without any contamination. In Figure 1, the $\hat{\epsilon}_{n}$ and $\hat{\delta}_{n}$ are plotted for the contaminated case and uncontaminated case in the 2nd and the 4th row respectively. Again, the max and average curves are computed for each combination of the constants $(C_{\phi}^{(1)},C_{\phi}^{(2)},C_{p})$ across multiple repetitions. We see a significant improvement through contamination, where $\hat{\epsilon}_{n}$ in the contaminated setup is only $1/10$ of the $\hat{\epsilon}_{n}$ in the uncontaminated setup while $\hat{\delta}_{n}$ curves are similar.
Remark 17 (Effect of Output Dimension).
We also ran the same setups but with parameter dimension $5$ and the results are plotted in Figure 2. We see that with lower dimensions, $\hat{\epsilon}_{n}$ are smaller in all cases. This is inevitable since a higher dimensional sample contains more information about the data set than a lower dimensional one.
Figure 2: Plots of estimated $(\epsilon_{n},\delta_{n})$ for the three setups with parameter dimension $5$.
(a) Linear regression case
(b) Logistic regression case
(c) Cauchy regression case

### v3

@@ 4.1 Estimation Procedure
In order for Assumption 4 to be satisfied for regression settings, one needs to impose a constraint on the size of covariates. In theory, we may allow the size of $\bm{W}$ to grow with $n$; for simplicity, we assume in all the simulations that $\|\bm{W}\|_{\infty}\leq 1$.
input : Data set size $n$; Contamination rate $p_{n}$; $\delta_{n}>0$; Likelihood function $f$; Contamination density $g$; prior $\pi_{0}$; Number of repeats $K$; Output quantile $q$; Number of particles $m$; True parameter $\theta^{*}$;
1 for $k=1,\dots,K$ do
2 if No dataset is present then
3      Generate true observations $\bm{X}:=\{\bm{x}_{1},\dots,\bm{x}_{n}\}$ according to $\bm{x}_{i}\sim f(\cdot;\theta^{*})$;
4
5   Contaminate the observations $\bm{X}$, denoted $\tilde{\bm{X}}$, by replacing each $\bm{x}_{i}$ with probability $p_{n}$ by a draw from $g(\cdot)$;
[…]
7   Sample $\theta_{1},\dots,\theta_{m}$ from the Laplace approximation;
[…]
9   Choose the minimal $\phi_{n}$ such that the total weight of particles outside $A_{n}:=B_{\phi_{n}}(\hat{\theta}^{*},\infty)$ is less than $\delta_{n}$  where $A_{n}$ is the $L_{\infty}$-ball around $\hat{\theta}^{*}$, i.e., a hypercube of width $2\phi_{n}$.;
@@ 11   Set $\mathcal{D}$ to be $\tilde{\bm{X}}$ with the last element discarded;
12   Compute the normalized weights $\tilde{w}_{i}$ for each $\theta_{i}$ (from Step 1) with respect to $\pi_{n-1}(\cdot;\mathcal{D})$;
@@ 13   Solve for
$\alpha:=\sup_{\bm{x}}\sum_{i=1}^{m}\tilde{w}_{i}d(\bm{x};\theta_{i}),\quad\beta:=\inf_{\bm{z}}\sum_{i=1}^{m}\tilde{w}_{i}d(\bm{x};\theta_{i})$
@@
14   Set $\hat{\epsilon}_{n}^{(k)}:=\log\eta_{n}+\log\alpha-\log\beta$;
15
@@ 14   Set $\hat{\epsilon}_{n}^{(k)}:=\log\eta_{n}+\log\alpha-\log\beta$;
17Sort $\epsilon_{n}^{(k)}$ in ascending order and output the $q$-th percentile.;
output : Estimated $\hat{\epsilon}_{n}$ with (100q)/100(100-q)/100 probability of failing ($\epsilon,\delta$)-DP.
@@
Algorithm 1 An empirical procedure to estimate $\epsilon$ given $\delta$ and model setup.
4.2 Posterior Sampling for Private Mean Estimation
In this section, we consider the private mean estimation problem, a setting that has been extensively investigated in the differential privacy literature, see, e.g., Biswas et al. (2020); Covington et al. (2021); Huang et al. (2021); Duchi et al. (2023), among others.
Let $\bm{X}:=\{x_{1},\dots,x_{n}\}$ denote a dataset of independent and identically distributed observations, where $x_{i}\sim\mathcal{N}_{[l,u]}(\theta^{*},\sigma^{2})$ are drawn from a truncated normal distribution supported on the interval $[l,u]$ with standard deviation $\sigma=8$. In the simulation study, the true mean is set to $\theta^{*}=30$, and the objective is to estimate $\theta^{*}$ under differential privacy constraints.
This subsection compares the performance of the following methods under an equivalent privacy budget:
  1. 1.
    Bayesian: draw $\hat{\theta}_{B}\sim\pi_{n}(\cdot|\bm{X})$;
  2. 2.
    CoinPress algorithm (Biswas et al., 2020);
  3. 3.
    Private Quantile Estimation + Clipped Mean Estimator (Huang et al., 2021);
For brevity, we omit the naive approach of injecting Laplace or Gaussian noise directly into the sample mean, as such methods perform significantly worse—by several orders of magnitude—under a conservative choice of interval size (ul=600u-l=600).
For the Bayesian contamination model, we consider a contamination rate of $p_{n}:=n^{-1/8}$ where $n$ is the sample size. The dataset $\bm{X}$ is contaminated by replacing each observation with probability $p_{n}$ by a draw from a scaled-and-translated Student T distribution $\tilde{x}_{i}\sim T_{\nu}(\theta^{*},\sigma)$ truncated to $[l,u]$. The density of $T_{\nu}(\theta^{*},\sigma)$ is given by
We employ a relatively diffuse prior $\pi_{0}(\cdot)\sim\mathcal{N}(40,40^{2})$ for $\theta$ and estimate $\hat{\epsilon}_{n}$ with fixed $\delta_{n}=\tfrac{1}{10n}$ using Algorithm 1. The estimated $\hat{\epsilon}_{n}$ values for varying $n$ are presented in Table 1. Sampling from the posterior is done through an ordinary random-walk Metropolis algorithm. We repeat the estimation for $1000$ simulated datasets, use the $99\%$ quantile of the resulting $\hat{\epsilon}_{n}$, and consider the Bayesian contamination mechanism to be $(\epsilon_{n},\delta_{n})$-DP with probability at least $99\%$.
The estimated $\hat{\epsilon}_{n}$ from the Bayesian approach is used as the privacy budget for the other two algorithms. However, both the CoinPress algorithm and the Clipped Mean Estimator provide privacy guarantees in zero-Concentrated DP (zCDP) (Dwork and Rothblum, 2016). The values $\hat{\epsilon}_{n}$ from Table 1 and $\delta_{n}=\tfrac{1}{10n}$ are used to solve for an appropriate $\rho_{n}$ budget for $\rho$-zCDP, by using the result that $\tfrac{\epsilon^{2}}{2}$-zCDP implies $(\epsilon\sqrt{\log(1/\delta)},\delta)$-DP, $\forall\delta>0$, (Bun and Steinke, 2016). Although this translation from $(\hat{\epsilon}_{n},\delta_{n})$-DP to $\rho_{n}$-zCDP imposes a stricter privacy restriction on frequentist approaches, it is a natural choice, e.g., the same translation would be adopted when comparing with the Gaussian mechanism under the $(\epsilon,\delta)$-DP framework.
Table 1: Estimated $\hat{\epsilon}_{n}$ values for different sample sizes $n$.
$n$ 100 1,000 2,000 5,000 10,000 20,000 50,000
$\hat{\epsilon}_{n}$ 2.85 0.94 0.72 0.46 0.32 0.26 0.18
To evaluate performance, we simulate a fresh dataset $\bm{X}$ and apply each of the three mechanisms to obtain estimates of $\theta^{*}$. For each dataset, the private mean estimation is repeated $1000$ times to compute the bias, variance, and mean squared error (MSE) of each estimator. This process is then repeated over $1000$ datasets to generate the box plots shown in Figure 1.
Figure 1 includes two versions of the CoinPress algorithm, which differ only by the number of iterations they take, $3$ and $10$, respectively. The same privacy budget is assigned to both versions. Intuitively, we see that when the number of iterations is not large enough, the performance of the estimator ceases to improve with $n$. Overall, a suitably tuned CoinPress algorithm outperforms other alternatives; however, the Bayesian approach achieves comparable performance to the Clipped Mean Estimator for $n\geq 1000$, with only a modest performance gap relative to CoinPress. It is worth emphasising that a posterior sample inherently contains richer information than a mean estimate. Also, we note that the variance of the MSE for the frequentist estimators is, in general, larger than that of the Bayesian estimator, which means that although the minimum MSE of the Bayesian estimator might not be as good as the frequentist methods, the performance is more consistent.
Remark 15.
While the three methods are evaluated on the same problem, their applicability differs. The CoinPress algorithm and the Clipped Mean algorithm are specifically tailored for private mean estimation, with the former assuming Gaussian-distributed data and the latter being model-free. Both essentially circumvent the need for a user-defined clipping bound, but this feature might not generalise to other problems. In contrast, the Bayesian approach naturally extends to a wider range of estimation tasks without altering the overall framework.
A potential drawback of introducing privacy through contamination is the absence of a straightforward, non-probabilistic relationship between the contamination rate and the resulting privacy level. By contrast, frequentist approaches, or more generally, methods that inject Gaussian or Laplace noise into a quantity with bounded sensitivity, offer an explicit correspondence between the privacy budget and the level of perturbation.
@@ 4.3 Simulation Setups and Results
Let $\bm{X}=\{X_{1},\dots,X_{n}\}\in\mathbb{R}^{n}$ be the set of observations and $\bm{W}\in[-1,1]^{n\times d}$ be the covariate matrix associated with $\bm{X}$, i.e., the $i$-th row of $\bm{W}$ is associated with the $i$-th observation $X_{i}$. By default, we assume the first column of $\bm{W}$ is a vector of $1$. Let $\theta\in\mathbb{R}^{d}$ denote the parameter, and the model likelihood is given by a density function $f(X_{i};\theta,\bm{w}_{i})=f(X_{i},\theta^{\top}\bm{w}_{i})$, such that $\int_{\mathbb{R}}f(x;\theta,\bm{w})\mathtt{d}x=1$ for any pair of $\theta,\bm{w}\in\mathbb{R}^{d}$. Let $\theta^{*}$ denote the true parameter value, which in practice may be replaced by an estimate $\hat{\theta}$, e.g., using the maximum a posteriori estimator.
For the cases where contamination is required, we choose a contamination density $g(x)$ and replace each $X_{i}$ with probability $p_{n}$ by a random sample from $g(\cdot)$. Note that the contamination density is allowed to depend on the covariates $\bm{w}$, see the example below for the linear regression case.
@@
$k_{p_{n}}(x;\theta,\bm{w})=(1-p_{n})f(x;\theta,\bm{w})+p_{n}\,g(x).$
The prior $\pi_{0}$ on $\theta$ is always set as a zero-mean Gaussian prior with uncorrelated dimensions and standard deviation $10$.
@@
    Bayesian Linear Regression:
    The regression model is given by $X_{i}=\theta^{*\top}\bm{w}_{i}+e_{i},$ where $e_{i}\sim\mathcal{N}(0,1)$. The likelihood function $f(X;\theta,\bm{w})$ is given by
    $f(x;\theta,\bm{w})=\frac{1}{\sqrt{2\pi}}\exp\left(-(x-\theta^{\top}\bm{w})^{2}/2\right).$
    The contamination density is chosen as $g(X_{i};\bm{w}_{i})\sim T_{5}(\theta^{*\top}\bm{w}_{i})$, a shifted Student’s t-distribution with degrees of freedom $5$, given by the following expression,
    $g(x)=\frac{\Gamma(3)}{\lambda\sqrt{3\pi}\Gamma(\tfrac{5}{2})}\left(1+\frac{x^{2}}{\lambda}\right)^{-3},$
    where $\lambda$ is the dispersion parameter, which may vary according to the dimension of $\theta$. In this setup, a heavy-tailed contamination is required in order to remove the compactness constraint on the observation space.
@@
    Bayesian Logistic Regression:
    The regression model is given by $X_{i}\sim\text{Ber}\left((1+e^{-\theta^{*\top}\bm{w}_{i}})^{-1}\right),$ where $X\in\{0,1\}$ with the likelihood function
    $f(x;\theta,\bm{w})=\left(1+e^{-\theta^{\top}\bm{w}_{i}}\right)^{-x}\left(1+e^{\theta^{\top}\bm{w}_{i}}\right)^{-1+x}.$
    Since the observation space is compact, $\epsilon$ is bounded in this setup. However, the simulation suggests that contamination should still be adopted. To contaminate the observations in this case, we replace each observation $X_{i}$ with probability $p_{n}$ by a draw from $\text{Ber}(\tfrac{1}{2})$. Thus, the contaminated density is $k_{p_{n}}(x;\theta,\bm{w})=(1-p_{n})f(x;\theta,\bm{w})+0.5p_{n}.$
@@
    Bayesian Cauchy Regression:
    The regression model is given by $X_{i}=\theta^{*\top}\bm{w}_{i}+e_{i}$ where $e_{i}\sim\text{Cauchy}(0)$ with the likelihood function
@@
    For contamination, we use another Cauchy distribution with a different scale
    $g(x)=\frac{1}{\lambda\pi\left[1+x^{2}/\lambda^{2}\right]}$
For each setting, we apply Algorithm 1 to estimate $\epsilon_{n}$ using simulated datasets with sizes ranging from 1,0001{,}000 to 500,000500{,}000 and record the $99$th percentile of the estimated $\epsilon_{n}$. In all simulations, $\delta_{n}$ is fixed to $(10n)^{-1}$ for numerical stability. The estimated quantiles of $\hat{\epsilon}_{n}$ for each setting are plotted in Figure 3 as solid lines with markers. The $\hat{\epsilon}_{n}$ are not log-scaled to better contrast the difference in magnitude between curves. We have included the same plots, but on a log scale, to better illustrate that $\hat{\epsilon}_{n}$ decreases polynomially with $n$, which is consistent with Theorem 4. We therefore fit the empirical decay rate of $\hat{\epsilon}_{n}$ and extrapolate the trend to larger sample sizes (up to $10^{7}$), shown as semi-transparent lines in the figure (without markers).
In the case of logistic regression, the likelihood ratio $d(x;\theta,\bm{w})$ is uniformly bounded $\forall x$ even without contamination. Nevertheless, Figure 2(b) shows that introducing contamination yields a substantial reduction in $\epsilon_{n}$, approximately by a factor of four.
The behaviour differs for Cauchy regression (Figure 2(c)), where the likelihood ratio $d(x;\theta,\bm{w})$ is likewise uniformly bounded $\forall x$. In this case, the uncontaminated setup achieves a lower estimated $\epsilon_{n}$, which is likely due to a faster contraction in the posterior distribution. Here, a smaller neighbourhood $A_{n}$ is shown to be more effective in controlling the likelihood ratio than the additional randomness introduced via contamination.
We also examine the effect of parameter dimensionality by estimating $\epsilon_{n}$ for both models with parameter dimensions $5$ and $10$. As expected, higher-dimensional parameter spaces incur a greater privacy cost, requiring larger sample sizes to achieve the same level of privacy. This issue is not intrinsic to our method but arises because the privacy leakage of an output is roughly proportional to its dimension, making this an intrinsically challenging setting. Our methodology does have the limitation that $\epsilon$ is difficult to control for small $n$, as the method achieves privacy through contamination more effectively when the contamination rate is moderate-to-low. This situation is analogous to applying insufficient noise to the output and trading privacy for statistical efficiency in the Laplace mechanism case. When the sensitivity scales linearly with the dimension of the function being privatised, achieving the same level of privacy requires the noise to also scale linearly with the dimension. This places our method close to a regime in which nearly the entire dataset must be contaminated in order to provide privacy.
Remark 16.
Differentially private Bayesian inference may also be achieved by introducing contamination at alternative stages of the inference procedure. Various such approaches have been proposed in the literature, e.g., Bernstein and Sheldon (2019); Kulkarni et al. (2021); Sheffet (2019). Although these methods provide non-probabilistic privacy guarantees, they typically rely on restrictive assumptions, such as bounded observation or parameter spaces, which our framework explicitly avoids. Moreover, the strength of their privacy guarantees is highly sensitive to the size of the imposed bounds and can deteriorate rapidly as these bounds increase. In contrast, our proposed framework is more general, as it contaminates the model at the data-level and does not require boundedness assumptions. Consequently, a fair comparison between these approaches is difficult, and we therefore omit a direct comparison.
Figure 2: Plots of estimated $\epsilon_{n}$ under the setting of Linear Regression, Logistic Regression and Cauchy Regression as the dataset size $n$ varies. The pale lines without nodes are extrapolated estimations.